# Label Leakage in Federated Inertial-based Human Activity Recognition

### v1

@@ 6. Discussion & Conclusions
In this paper, we presented a first-of-its-kind analysis of label reconstruction in federated HAR. We conducted a comprehensive evaluation of five gradient inversion techniques, one of which was introduced in this work, and demonstrated that the unbalanced nature of HAR datasets poses significant privacy risks that can be exploited by leakage attacks. We further showed that LDP measures, namely gradient clipping and noise addition, are partially effective when combined, but insufficient to ensure full privacy in cases involving sequential sampling. Regardless of whether leakage methods were applied to multi-step averaged gradients or LDP-distorted gradients, label-exclusive batches still allowed certain attacks to infer the presence of classes with high accuracy. In general, our analysis suggests that the number of activity classes, the degree of class imbalance, and the sampling strategy are the main factors influencing the extent of label leakage.

### v2

@@ 6. Discussion & Conclusions
This paper presented a first-of-its-kind analysis of label reconstruction in federated HAR. A comprehensive evaluation of five gradient inversion techniques demonstrated that the unbalanced nature of HAR datasets poses significant privacy risks that can be exploited by leakage attacks. LDP measures, namely gradient clipping and noise addition, proofed to be partially effective when combined, but insufficient to ensure full privacy in cases involving sequential sampling. Regardless of whether leakage methods were applied to multi-step averaged gradients or LDP-distorted gradients, label-exclusive batches still allowed certain attacks to infer the presence of classes with high accuracy. Our analysis suggests that the number of activity classes, the degree of class imbalance, and the sampling strategy are the main factors influencing the extent of label leakage.